# Sample Complexity of Composite Quantum Hypothesis Testing

### v1

@@ I Introduction
Building on these recent developments, this paper provides the first comprehensive characterization of the sample complexity of symmetric composite QHT. Our key contributions are as follows: First, we derive new lower and upper bounds on the sample complexity for various classes of uncertainty sets, including both finite and infinite cardinalities. Specifically, for testing a singleton pure uncertainty set $\mathcal{D}_{1}=\{\left|\psi\right\rangle\!\left\langle\psi\right|\}$ against a composite uncertainty set $\mathcal{D}_{2}$, we show that the sample complexity scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(\frac{1}{\delta})}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\Bigr)$. We generalize this result to the case when both the sets are finite, showing that the sample complexity is governed by the maximum pairwise fidelity between the sets, $F_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$, as $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|/\delta)}{-\ln F_{\max}}\Bigr)$. We also extend this analysis to infinite cardinality uncertainty sets under specific fidelity constraints. Secondly, we extend our framework to the locally differentially private setting, establishing the first sample complexity bounds for private composite QHT.
@@ II-B Simple Quantum Hypothesis Testing
In simple binary QHT, a quantum system is prepared in one of two known states $\rho_{1}^{\otimes}$ or $\rho_{2}^{\otimes n}$. The distinguisher is given access to $n$ identical copies of the state without a priori knowledge of which was prepared. The objective of the distinguisher is to determine the identity of the state by designing a two-outcome positive operator-valued measurement (POVM) $\{\Pi,I-\Pi\}$, where the outcome associated with the operator $0\leq\Pi\leq I$ corresponds to guessing $\rho_{1}^{\otimes n}$ and the outcome associated with $I-\Pi$ corresponds to guessing $\rho_{2}^{\otimes n}$.
@@ III-C2 Uncertainty Sets of Finite Cardinality
Suppose that $|\mathcal{D}_{i}|=m_{i}<\infty$. Then,
@@
$=\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
@@
Furthermore, for fixed $p$, (14) and (17) together imply that $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(m_{1}m_{2}/\delta)}{-\ln F_{\max}}\Bigr).$
@@ A CPTP map $\mathcal{M}$ is $\varepsilon$-LDPQ for $\varepsilon\geq 0$ if
Let $\mathcal{M}$ be an $\varepsilon$-LDPQ channel applied independently to each of the $n$-copies of the unknown quantum state. We then denote the locally private uncertainty sets and their corresponding convex hulls respectively as
$\mathcal{D}_{i,n}^{\mathcal{M}}:=\{\mathcal{M}(\rho_{i})^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},\quad\mathcal{C}_{i,n}^{\mathcal{M}}:=\operatorname{conv}(\mathcal{D}_{i,n}^{\mathcal{M}}),$
@@
Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$. Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(e^{\varepsilon}+1)\ln\frac{p(1-p)}{\delta}}{2(e^{\varepsilon}-1)H_{1/2}(\rho_{1}\|\rho_{2})}$. Additionally, [7, (4.25)] yields that $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(1-\delta/p)^{2}}{e^{-\varepsilon}(e^{\varepsilon}-1)^{2}E_{1}(\rho_{1}\|\rho_{2})^{2}}$. The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$. ∎
The above result directly generalizes the sample complexity of differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
@@
For the setting of Proposition˜15, we have
$\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}\leq-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max},$ (24)
@@
$-\ln\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (25)
$\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (26)
$\geq\frac{1}{2}\sup_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2}))^{2}$ (27)
$\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$ (28)
where in (26), we use that $1-x\leq-\ln x$ for all $x\in[0,1]$. (27) uses [7, Equation 4.7], and (28) uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$. ∎

### v2

@@ I Introduction
Building on these recent developments, this paper provides the first comprehensive characterization of the sample complexity of symmetric composite QHT. Our key contributions are as follows: First, we derive new lower and upper bounds on the sample complexity for various classes of uncertainty sets, including both finite and infinite cardinalities. Specifically, for testing a singleton pure uncertainty set $\mathcal{D}_{1}=\{\left|\psi\right\rangle\!\left\langle\psi\right|\}$ against a composite uncertainty set $\mathcal{D}_{2}$, we show that the sample complexity scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(\frac{1}{\delta})}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\Bigr)$. We generalize this result to the case when both the sets are finite, showing that the sample complexity is governed by the maximum pairwise fidelity between the sets, $F_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$, as $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(\sqrt{|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|/}\delta)}{-\ln F_{\max}}\Bigr)$. We also extend this analysis to infinite cardinality uncertainty sets under specific fidelity constraints. Secondly, we extend our framework to the locally differentially private setting, establishing the first sample complexity bounds for private composite QHT.
@@ II-B Simple Quantum Hypothesis Testing
In simple binary QHT, a quantum system is prepared in one of two known states $\rho_{1}^{\otimes n}$ or $\rho_{2}^{\otimes n}$. The distinguisher is given access to $n$ identical copies of the state without a priori knowledge of which was prepared. The objective of the distinguisher is to determine the identity of the state by designing a two-outcome positive operator-valued measurement (POVM) $\{\Pi,I-\Pi\}$, where the outcome associated with the operator $0\leq\Pi\leq I$ corresponds to guessing $\rho_{1}^{\otimes n}$ and the outcome associated with $I-\Pi$ corresponds to guessing $\rho_{2}^{\otimes n}$.
@@ III-C2 Uncertainty Sets of Finite Cardinality
Let $|\mathcal{D}_{i}|=m_{i}<\infty$ for $i=1,2$. Then,
@@
$\leq\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
@@
Furthermore, for fixed $p$, (14) and (17) together imply that $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(\sqrt{m_{1}m_{2}}/\delta)}{-\ln F_{\max}}\Bigr).$
@@ A CPTP map $\mathcal{M}$ is $\varepsilon$-LDPQ for $\varepsilon\geq 0$ if
Let $\mathcal{M}$ be an $\varepsilon$-LDPQ channel applied independently to each of the $n$-copies of the unknown quantum state. We then denote the locally private uncertainty sets as
$\mathcal{D}_{i,n}^{\mathcal{M}}:=\{\mathcal{M}(\rho_{i})^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},$
@@
Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$. Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2). Additionally, [7, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2. The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$. ∎
The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
@@
Consider the setting of Proposition˜15. Then, the following inequality holds,
$\frac{1}{2}\biggl(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\biggr)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}\leq-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max},$ (24)
@@
$-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max}$ $\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$
$\geq\frac{1}{2}\sup_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2}))^{2}$
$\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$
where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [7, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$. ∎